OpenStudy (anonymous):

what is the measure of side cb if ba is 17 and ca is 15

OpenStudy (jdoe0001):

having the picture posted would help

OpenStudy (anonymous):

sorry its a right triangle

OpenStudy (anonymous):

use the a^2+b^2=c^2

OpenStudy (anonymous):

which side is missing one of the legs or the hypotenues?

OpenStudy (anonymous):

side cb

OpenStudy (mrnood):

is cb the longest side? (opposite the right angle? or one of the other sides?

OpenStudy (anonymous):

is side cb the longest side?

OpenStudy (anonymous):

ba which is 17 is the hypotenuse

OpenStudy (anonymous):

ok so then subtract 17^2 from 15^2

OpenStudy (anonymous):

and ca which is 15 is the bottom

OpenStudy (mrnood):

 x^2 +15^ = 17^2

OpenStudy (jdoe0001):

got picture?

OpenStudy (anonymous):

yes that's it ^^

OpenStudy (jdoe0001):

 \(\bf c^2={\color{blue}{ a}}^2+b^2\implies c^2-b^2={\color{blue}{ a}}^2\implies \sqrt{c^2-b^2}={\color{blue}{ a}}\\ \qquad \uparrow \\ \textit{pythagorean theorem}\)